OpenStudy (anonymous):

Given the function f(x) = −2x + 8, find x if f(x) = 22.

OpenStudy (ash2326):

we have \[f(x)=-2x+8\] f(x)=22 \[22=-2x+8\] find x from here, can you do this?

OpenStudy (anonymous):

i got -19

OpenStudy (ash2326):

\[22=-2x+8\] \[22-8=-2x\] \[14=-2x\] so \[x=?\]

OpenStudy (anonymous):

i though you did multiplication first?

OpenStudy (ash2326):

where did I do?

OpenStudy (ash2326):

I think you did \[-2(x+8)\] but it's \[-(2x)+8\] so \[22=-2x+8 \]

OpenStudy (anonymous):

you did subtraction. But now i believe that was right

OpenStudy (ash2326):

yeah:) do you follow @vetdoc ?

OpenStudy (anonymous):

yes... :P i get it now

OpenStudy (ash2326):

great:D